Math:50:640:041 Elementary Algebra Summer 2010

Course Specifics:

Instructor: Tejas Gandhi

E-mail:  tejasg@camden.rutgers.edu

Classroom: ATG 123 Office Hours: by appointment

If you have any questions and cannot make it to my office hours please send me email and we can arrange to meet.

Text Book:

The required text for this course is the following. Elementary and Intermediate Algebra, 5^{th} edition, Jerome Kaufman.

Course Objectives:

The basic requirement upon completion of this couses, is the student’s ability to utilize and apply fundamental math concepts. The student will demonstrate a proficiency with the operations with integers, rational expressions, polynomials, expressions involving exponents and radicals, solving one variable equations, graphing lines, solving literal equations and related word problems.

General Syllabus:

The following chapters/sections will be covered: 1, 2, 3, 4.1-4.4, 5.1-5.4, 6, 7.1-7.4, 9.1-9.5, 10.1-10.4.

Grading:

Your final grade in the course will be decided by your performance in 4 exams.

1. Exam 1 – 25%.

2. Exam 2 – 25%.

3. Exam 3 – 25%.

4. Comprehensive final – 25%.

50:640:042 2

The final exam is mandatory (required by the State). Missed exams cannot be made up unless arrangements are made prior to the exam date.

Attendance:

There are no excused absences. Attendance of all classes is expected. Any material or announcements missed is the full responsibility of the student.